OpenStudy (anonymous):

So this question had me stumped on my test earlier. Find the surface area: y=sqrt(1-(x^2/4)) where 0 is less than or equal to x less than or equal to 2

OpenStudy (anonymous):

\[y=\sqrt{1-x^2/4}\]

OpenStudy (anonymous):

\[0\le x \le2\]

OpenStudy (turingtest):

surface area? is this a revolution?

OpenStudy (anonymous):

And the formula is \[\int\limits_{a}^{b}2\Pi yds\]

OpenStudy (anonymous):

\[ds=\sqrt{1+(dy/dx)^2}\]

OpenStudy (turingtest):

so you are revolving around the x-axis?

OpenStudy (anonymous):

Yesh, sorry for not answering

OpenStudy (anonymous):

about the x-axis

OpenStudy (turingtest):

Ok here's an idea...

OpenStudy (anonymous):

Any ideas are much appreciated!

OpenStudy (turingtest):

\[y=\sqrt{1-\frac{x^2}4}\implies y^2=1-\frac{x^2}4\]using implicit differentiation we get\[2yy'=-\frac x2\implies y'=-{x\over4y}\]\[2\pi\int yds=2\pi\int y\sqrt{1+{x^2\over16y^2}}dx=2\pi\int_{0}^{2}\sqrt{y^2+{x^2\over 16}}dx\]I'll continue, but let you digest in the meantime...

OpenStudy (anonymous):

Sounds good, willing to wait

OpenStudy (anonymous):

Wait, is not ok to leave it in terms of x since we are rotating about the x-axis?? and by "y" in the formula, it means whatever y=. Just to clear up some confusion.

OpenStudy (turingtest):

but this is where the trick of implicit differentiation pays off: we will now sub in that expression for y into the formula above:\[2\pi\int_{0}^{2}\sqrt{1-{x^2\over4}+{x^2\over16}}dx=2\pi\int_{0}^{2}\sqrt{1-{3x^2\over16}}dx\]and now we have a trig sub... I hope

OpenStudy (turingtest):

more specifically we subbed in for \(y^2\)

OpenStudy (anonymous):

Haha oh ok. I was thinking I would need a trig substitution, but radical was not very pretty looking

OpenStudy (turingtest):

yeah, but waiting to sub in for y made it a lot easier

OpenStudy (anonymous):

cause my radical looked very similar but my trig substitution was not making any sense. This is what i had on my test, which I know is wrong. \[\pi/2\int\limits_{0}^{2} \sqrt{-3x^2+16}\]

OpenStudy (turingtest):

it is the same as mine, but with pi/2 for some reason I don't get...

OpenStudy (anonymous):

Well I had 16 on bottom, so i split the radical up into to with numerator squared over the denominator squared. And the sqrt(16) is 4 so I took out the constant of 1/4 and that times 2pi=pi/2. But my trig sub. was not working out to nicely, so I figured I messed up somewhere.

OpenStudy (turingtest):

oh yeah, my algebra was off you are right about the pi/2 that said, we have the same thing\[\sqrt{-3x^2+16}=\sqrt{16-3x^3}=\frac14\sqrt{1-{3x^2\over16}}\]so try the sub\[\frac4{\sqrt3}\sin\theta=x\]

OpenStudy (anonymous):

ohhhh that is where I messed up, all i did was \[1/\sqrt{3}\] \[\sin \theta\]

OpenStudy (anonymous):

well those are suppose to be together. haha

OpenStudy (turingtest):

gotta get rid of that 16 ;) I gotta go, glad I could help!

OpenStudy (anonymous):

Yes, thank you very much!!!